# Confidence-driven Sampling for Backdoor Attacks

### v1

@@ 5.3 Performance of CBS in Type II backdoor attacks
Model ResNet18 ResNet18 normal-$\rightarrow$ VGG16
@@
Performance comparison. As presented in Table 2, our method CBS, demonstrates similar behavior to Type I attacks, showing enhanced resistance against various defense mechanisms at the cost of some success rate. Notably, CBS consistently outperforms Random and FUS when defenses are in place, highlighting its versatility in different scenarios. Particularly for vulnerable attacks like BadNet, CBS achieves substantial improvements, surpassing Random by over $10\%$ and FUS by over $5\%$. Furthermore, CBS exhibits smaller standard errors, indicating its higher stability. However, there is still room for further improvement, as the absolute success rate is not as impressive as in Type I attacks. We consider this area for future research endeavors.
5.4 Performance of CBS in Type III backdoor attacks
Attacks & Defenses. We consider 3 Representative attacks in this category—Lira (Doan et al., 2021b) which involves a stealthy backdoor transformation function and iteratively updates triggers and model parameters; WaNet (Nguyen & Tran, 2021) which applies the image warping technique to make triggers more stealthy; Wasserstein Backdoor (WB) (Doan et al., 2021a) which directly minimizes the distance between poisoned and clean representations. Note that Type III attacks allow the attackers to take control of the training process. Though our threat model does not require this additional capability of attackers, we follow this assumption when implementing these attacks. Therefore, we directly select samples based on ResNet18 and VGG16 rather than using ResNet18 as a surrogate model. We conduct 3 representative defenses that are applicable for this type of attacks—Neural Cleanser (NC) (Wang et al., 2019), STRIP (Gao et al., 2019), Fine Pruning (FP) (Liu et al., 2018). We follow the default settings to implement these attacks and defenses (details in Appendix A.2). We set $\epsilon=0.37$ which matches the poison rate $p=0.1$ in the original settings of backbone attacks. Results on Cifar10 and Cifar100 are presented in Table 3.
Performance comparison. Except for the common findings in previous attacks, where CBS consistently outperforms baseline methods in nearly all experiment, we observe that the impact of CBS varies when applied to different backbone attacks. Specifically, CBS tends to yield the most significant improvements when applied to WB, while its effect is less pronounced when applied to WaNet. For example, when confronting FP and comparing CBS with both Random and FUS, we observed an increase in ASR of over $7\%$ on WB, while the increase on WaNet amounted to only $3\%$, with Lira showing intermediate results. This divergence may be attributed to the distinct techniques employed by these attacks to enhance their resistance against defenses. WB focuses on minimizing the distance between poisoned samples and clean samples from the target class in the latent space. By selecting boundary samples that are closer to the target class, WB can reach a smaller loss than that optimized on random samples, resulting in improved resistance. The utilization of the fine-tuning process and additional information from victim models in Lira enable a more precise estimation of decision boundaries and the identification of boundary samples. WaNet introduces Gaussian noise to some randomly selected trigger samples throughout the poisoned dataset, which may destroy the impact of CBS if some boundary samples move away from the boundary after adding noise. These observations suggest that combining CBS with proper trigger designs can achieve even better performance, and it is an interesting topic to optimize trigger designs and sampling methods at the same time for more stealthiness, which leaves for future exploration.

### v2

@@ 5.2 Performance of CBS in Type I backdoor attacks
Model ResNet18 ResNet18 $\rightarrow$ VGG16
@@
Performance comparison. Generally, CBS enhances the resilience of backbone attacks against various defense mechanisms. It achieves notable improvement compared to Random and FUS without a significant decrease in ASR when no defenses are in place. This is consistent with our analysis in Section 4.3. We notice that CBS has the lowest success rate when no defenses are active, which is consistent with our analysis in Remark 4.8141414Additional discussion can be found in Appendix 8.8. Nonetheless, CBS  still achieves commendable performance, with success rates exceeding $70\%$ and even reaching $90\%$ for certain attacks. It is important to note that the effectiveness of CBS varies for different attacks and defenses. The improvements are more pronounced when dealing with stronger defenses and more vulnerable attacks. For instance, when facing SS, a robust defense strategy, CBS significantly enhances ASR for nearly all backbone attacks, especially for BadNet. In this case, CBS can achieve more than a $20\%$ increase compared to Random and a $15\%$ increase compared to FUS. Additionally, it’s worth mentioning that CBS consistently strengthens resistance against detection-based (first two) and non-detection-based defenses (the other two). This further supports the notion that boundary samples are inherently more challenging to detect and counteract. While the improvement of CBS on VGG16 is slightly less pronounced than on ResNet18, it still outperforms Random and FUS in nearly every experiment, indicating that CBS can be effective even on unknown models.
5.3 Performance of CBS in Type II backdoor attacks
Attacks & Defenses. We consider 2 representative attacks in this category—Hidden-trigger (Saha et al., 2020), which adds imperceptible perturbations to samples to inject backdoors, and Clean-label (LC) (Turner et al., 2019), which leverages adversarial examples to train a backdoored model. We follow the default settings in the original papers and adapt $l_{2}$-norm bounded perturbation (perturbation size $6/255$) for LC. We test all attacks against three representative defenses that are applicable to these attacks. We include NC, SS, Fine Pruning (FP) (Liu et al., 2018), Anti-Backdoor Learning (ABL) (Li et al., 2021a). We set $\epsilon=0.3$ for CBS and $p=0.2\%$ for Random and FUS correspondingly. For every experiment, a source class and a target class are randomly chosen, and poisoned samples are selected from the source class. The success rate is defined as the probability of misclassifying samples from the source class with triggers as the target class. Results on dataset Cifar10 and Cifar100 are presented in Table 2. We include additional results on Tiny-ImageNet in the Supplementary for a further illustration.
@@ 5.3 Performance of CBS in Type II backdoor attacks
Performance comparison. As detailed in Table 2, CBS displays an enhanced capacity to withstand various defense mechanisms, akin to Type I attacks, while sacrificing a marginal degree of success rate. Notably, in the presence of defenses, CBS consistently surpasses both Random and FUS strategies in performance, demonstrating its adaptability across various challenging conditions. This is particularly evident when tackling susceptible attack strategies like BadNet, where CBS not only achieves substantial gains—outperforming Random by upwards of $10\%$ and FUS by more than $5\%$—but also maintains smaller standard errors. These smaller errors reflect CBS’s stability, which is critical in real-world applications where consistent performance is crucial.
5.4 Performance of CBS in Type III backdoor attacks
Attacks & Defenses. We consider 3 Representative attacks in this category—Lira (Doan et al., 2021b) which involves a stealthy backdoor transformation function and iteratively updates triggers and model parameters; WaNet (Nguyen & Tran, 2021) which applies the image warping technique to make triggers more stealthy; Wasserstein Backdoor (WB) (Doan et al., 2021a) which directly minimizes the distance between poisoned and clean representations. Note that Type III attacks allow the attackers to take control of the training process. Though our threat model does not require this additional capability of attackers, we follow this assumption when implementing these attacks. Therefore, we directly select samples based on ResNet18 and VGG16 rather than using ResNet18 as a surrogate model. We conduct 3 representative defenses that are applicable for this type of attacks—NC, STRIP, FP. We follow the default settings to implement these attacks and defenses. We set $\epsilon=0.37$ which matches the poison rate $p=0.1$ in the original settings of backbone attacks. Results on Cifar10 and Cifar100 are presented in Table 3.
Performance comparison. Except for the common findings in previous attacks, where CBS consistently outperforms baseline methods in nearly all experiments, we observe that the impact of CBS varies when applied to different backbone attacks. Specifically, CBS tends to yield the most significant improvements when applied to WB, while its effect is less pronounced when applied to WaNet. For example, when confronting FP and comparing CBS with both Random and FUS, we observed an increase in ASR of over $7\%$ on WB. In comparison, the increase on WaNet amounted to only $3\%$, with Lira showing intermediate results. This divergence may be attributed to the distinct techniques employed by these attacks to enhance their resistance against defenses. WB focuses on minimizing the distance between poisoned samples and clean samples from the target class in the latent space. By selecting boundary samples that are closer to the target class, WB can reach a smaller loss than that optimized on random samples, resulting in improved resistance. The utilization of the fine-tuning process and additional information from victim models in Lira enable a more precise estimation of decision boundaries and the identification of boundary samples. WaNet introduces Gaussian noise to some randomly selected trigger samples throughout the poisoned dataset, which may destroy the impact of CBS if some boundary samples move away from the boundary after adding noise. These observations suggest that combining CBS with proper trigger designs can achieve even better performance, and it is an interesting topic to optimize trigger designs and sampling methods at the same time for more stealthiness, which leaves for future exploration.
5.5 CBS with Partial Backdoor
We investigate scenarios where attackers can only manipulate partial training data. Specifically, we conduct experiments on the ResNet18 model using the Cifar10 dataset, employing the BadNet attack method with various sampling strategies. We designate different subset rates (10%, 5%, 1%) of the training set as accessible to the attacker, who can only poison this fraction of the data. From their accessible data, attackers insert triggers into 10% of the samples. The effectiveness of these attacks, under different defense mechanisms, is evaluated. Our findings, presented in Table 4, demonstrate that our method effectively enhances the stealthiness of backdoor attacks, even with limited data access. This underscores the practical potential of our approach in real-world situations where attackers cannot access the entire training dataset.
Figure 6: Illustrating impacts of confidence.
Table 4: Experiments for partially poisoned data. Conducted on model ResNet18 and dataset Cifar10, attacking method BadNet is incorporated with different sampling methods.
Subset rate Random FUS CBS
No defenses 10% 99.9 99.9 93.7
5% 98.4 99.7 92.9
1% 97.2 99.7 92.3
SS 10% 1.2 6.8 15.7
5% 0.9 5.3 12.4
1% 0.6 2.8 8.5
NC 10% 2.7 8.2 14.4
5% 1.4 6.5 10.7
1% 2.2 5.3 8.4
Strip 10% 1.5 4.9 13.2
5% 0.9 3.1 9.5
1% 0.5 2.8 7.2
@@ 5.6 Ablation study
We also conduct experiments to study the effect of poisoning rate, and due to the page limit, we include it in Appendix 8.9.